OpenStudy (anonymous):

suppose that (g(x))^2+6x=x^2g(x)+17 and g(2) =5 find g'(2)

OpenStudy (anonymous):

Please.. i need help for the answer.. this is due tonight and i need to get it done!

OpenStudy (freckles):

Have you differentiate the right side yet? That is the side I seen you last working on.

OpenStudy (anonymous):

yes, 3g(x)2?

OpenStudy (anonymous):

3g(x)^2

OpenStudy (freckles):

wait are you trying to do left side or tight side because well that is neither side differentiated correctly...

OpenStudy (freckles):

\[\frac{d}{dx}(g(x))^2=2g(x)\cdot g'(x) \ \text{ By chain rule } \\ \frac{d}{dx}6x =?\text{ I will leave this one for you } \\ \frac{d}{dx} (x^2g(x))=(x^2)'g(x)+x^2(g(x))' \text{ By product rule } \]

OpenStudy (anonymous):

6

OpenStudy (freckles):

ok and finish up that last line too

OpenStudy (anonymous):

i was trying to do the right side, isnt the left side 2(g(x))g'(x)+6?

OpenStudy (freckles):

yes your left hand side is right

OpenStudy (anonymous):

ok then is the right side 2x*g(x)

OpenStudy (anonymous):

the g(X) throws me off a little

OpenStudy (freckles):

did you see my last line above with the pretty latex?

OpenStudy (anonymous):

ok so its (x^2) g(x) +x^2(g(x))

OpenStudy (freckles):

You need to differentiate the parts that are in ( )'

OpenStudy (anonymous):

x^2' g(x) + x^2 (g(x))'

OpenStudy (freckles):

(x^2)'=2x... (g(x))'=g'(x)

OpenStudy (anonymous):

would 2x stay prime or no

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

i got confused for a sec

OpenStudy (freckles):

(x^2)''=(2x)'=2 But we had (x^2)'

OpenStudy (freckles):

and (x^2)'=2x

OpenStudy (anonymous):

2x g(x) +x^2 g(x)

OpenStudy (freckles):

not exactly

OpenStudy (freckles):

Are you not familiar with product rule?

OpenStudy (anonymous):

i did learn it but i was last week so its not clear in my head

OpenStudy (anonymous):

it*

OpenStudy (freckles):

\[(uv)'=u'v+v'u \\ (x^2g(x))'=(x^2)'g(x)+g'(x)x^2 \\ (x^2g(x))'=2xg(x)+g'(x)x^2 \]

OpenStudy (anonymous):

okay thank you, now do i plug in 2?

OpenStudy (freckles):

yes replace the x's with 2

OpenStudy (freckles):

then solve for g'(2)

OpenStudy (freckles):

don't forget to replace g(2) with 5

OpenStudy (anonymous):

ok.. to verify my answeer

OpenStudy (anonymous):

freckles

OpenStudy (freckles):

\[2g(x)g'(x)+6=2xg(x)+g'(x)x^2 \\ 2g(x)g'(x)-g'(x)x^2=2xg(x)-6 \\ g'(x)[2g(x)-x^2]=2xg(x)-6 \\ g'(x)=\frac{2xg(x)-6}{2g(x)-x^2}\]

OpenStudy (freckles):

Replace x's with 2's

OpenStudy (freckles):

\[g'(2)=\frac{2(2)g(2)-6}{2g(2)-2^2}\]

OpenStudy (freckles):

you are give g(2) is 5 so replace g(2) with 5

OpenStudy (anonymous):

yeah, did u get 14/6

OpenStudy (anonymous):

yup its right